From: STOSC::SMM "On-Duty SMM x5074" 26-JAN-1993 23:25:02.82 To: @[SMM.DIS]HST_STAT CC: Subj: HST STATUS REPORT, 01/25 - 01/26 HUBBLE SPACE TELESCOPE DAILY REPORT #809 PERIOD COVERED: 0800 EST 01/25/93 - 0800 EST 01/26/93 Daily Status Report as of 026/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOC/96 4152 (Observations to Monitor Desorption) The Faint Object Camera (f/96) was used to take a series of four images using standard star GRW58211 to monitor the focus of the Optical Telescope Assembly (OTA). All observations were executed in fine lock. A corrective SIP maneuver was successfully uplinked. All four images were reported to be well centered. 1.2 Completed FOC/96 4069 (Far-UV Spectra of High Red-Shift Quasars) The Faint Object Camera was used to take three ultraviolet objective prism spectral images of the high red shift quasar Q0120+027. All three images were taken in fine lock. A couple of short recentering episodes were noted during the real-time telemetry on the third image. All images were reported to look okay. 1.3 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take four high latitude survey images which were taken in parallel with the prime science observations. There were no problems, except a couple of short episodes of recentering. The images were received and displayed. They were reported to look okay. 1.4 Completed FOS/RD 3840 (Evolution of Carbon, Nitrogen and Oxygen in ST) The Faint Object Spectrograph (Red Side) was used to take three ultraviolet spectral observations of the extra galactic H II regions around galaxy C1543+091. The binary search was successful, and all three of the science observations were reported to be satisfactory. 1.5 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.6 Completed WFC 4105 (HST Medium-Deep Survey: High Latitude) The Wide Field Camera was used to take two medium-deep sky survey images which were taken in parallel with the primary science observations. Both images were reported to be satisfactory. 1.7 Completed FOS/RD 4698 (FOS Dark Monitoring) The Faint Object Spectrograph was used on both the red and blue sides to obtain two dark exposures to determine the instruments background calibration. The first observation was reported to show several spikes with count levels at about 200, while the second observation only had one or two counts per diode. 1.8 Completed GHRS/2 3444 (Gas in Interstellar Clouds) The High Resolution spectrograph was used to make spectral observations of early-type star HD36486 looking for interstellar gas clouds. The observations were executed as scheduled, and no problems were reported. 1.9 Start of Zero-Gyro (On-Orbit) Test 026/1200Z through 028/0000Z. (See note 2.4) . 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 8 Successful : 8 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Note: The Zero-Gyro (On-Orbit) Test was successfully initiated today at 026/1200Z, and is scheduled to last through Thursday 028/0000Z. No operational problems have been reported. The Fight Readiness Review meeting (FRR) was held last Friday, Jan. 22, and a final review was held yesterday. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Continuation of the Zero-Gyro (On-Orbit) Test from 026/1200Z through 028/0000Z. 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-